A hot air balloon has a volume of 20 cubic meters. If the air inside is at a temperature of 70 degrees C and the outside air has a density of 1.17 kg/cubic meter, how much lift does the balloon generate? Hint: use 29.0 grams/mole of air.
Answer #1

V = 20 m^3

T = 273 + 70 = 343 K

P V = n R T

(101325) (20) = n (8.314) (343)

n = 710.6 mol

mass of gas inside = 29 x 710.6 = 20.6 x 10^3 grams

= 20.6 kg


Buoyant force, Fb = (20x 1.17)(9.8) = 229.32 N upward  

left by balloon = (229.32) - (20.6 x 9.8)

= 27.44 N .........Ans